# ”Living Within Four Walls”: Exploring Emotional and Social Dynamics in Mobile Usage During Home Confinement

### v2

@@ 5.2. Unsupervised Activity Recognition From App Sequences
Segmentation. To study how the application usage is related to the aforementioned variables, we propose App-based Information Gain Temporal Segmentation (AIGTS), which focuses on differentiating activities within app usage. Figure 6 provides an illustration of how activities are extracted from app usage data. AIGTS draws inspiration from Information Gain Temporal Segmentation (IGTS) (80), an unsupervised method for segmenting multivariate time series data. IGTS works by identifying which splits in the multivariate time series minimize the weighted entropy across the series. The variables within the multivariate time series data corresponded to whether an app of a particular category was used.
@@
Unlike IGTS focusing solely on the time series, AIGTS aims to identify different segments and transition times between apps throughout the day. It is capable of detecting when participants change their app usage patterns, such as transitioning from using Email and Browsers (both categorized under Communication) to browsing Reddit (News and Media category). This method allows for segmentation at any level of granularity desired, ranging from half a day to as little as half a minute for application activity. The granularity of segmentation is determined by the variable $k$, which represents the number of splits made in the time series. To ensure that segments are captured at a consistent level of granularity, we propose the following steps for calculating the optimal $k$ in AIGTS:
  1. (1)
    Calculate the knee point of the activity data for each day;
  2. (2)
    Identify the average number of segments per minute for all days based on the knee point;
  3. (3)
    Compute the median average segments per minute for all users;
  4. (4)
    Multiply this median by the number of minutes per day (which varies from the first log to the last log) and choose this value as $k$.
These procedures guarantee that the average length of each segment will be consistent across all days.
@@
Figure 6. An illustration of unsupervised activity recognition using app usage data
@@
Clustering. Next, AIGTS was conducted on all-time series, resulting in a segmented list for each user. To extract features from these segments, k-means clustering (51) was applied. The distribution of different categories appearing in a segment was treated as an $n$ dimensional vector, where $n$ is the number of app categories. To identify the optimal number of segments, Silhouette analysis (52) was utilized with a maximum of 20 clusters. Especially, the label indicating which cluster a segment belongs to was used as a feature representing the activity within that segment. These segment labels were derived manually from the average distribution of apps in that cluster. In this paper, a segment that belongs to a particular cluster will be referred to as an ‘activity’.
Distribution of User Activities. Following the segmentation and clustering, we uncovered 16 clusters. The average duration of these user activities was 18 minutes. Figure 7 displays average activities for all clusters, where the average activity represents the average distribution of app categories appearing in each segment within the cluster. These clusters were manually named to describe the distribution of average activities. The $x$-axis displays the app categories that frequently appear in clusters, with infrequently appearing categories omitted. Notably, the cluster Other includes all uncategorizable segments and has roughly equal means across multiple categories. Additionally, the cluster Idling is used for Personalization apps because Personalization apps are typically launchers (see table 2), and indicate that the user is sitting on their device’s home screen. Moreover, the most frequent user activities during the COVID-19 lockdown were Just Communication and Idling, accounting for 17% and 14.60% of activities, respectively. Conversely, the least frequent activities were Business Communication and Music, accounting for 1.22% and 1.65% of activities, respectively.
Figure 7. Average activities for all clusters
5.3. App Usage Correlation
Furthermore, we investigate the correlations between user activities with their valence, arousal, productivity, and interruptions. To understand whether any particular activity $A$ was correlated with valence, we ran a Spearman correlation (24) between valence scores and the amount of time spent in user activity $A$ in the previous hour. The same goes for the arousal, productivity, and interruptions, but the productivity and interruptions vary in duration from 1 hour to 24 hours, as these are calculated from the end-of-day survey and correspond to the value for a full day. Although the ESM responses also measure interruptibility, we used the interruptibility item from the end-of-day survey to assess the intensity of how the COVID lockdown affected well-being, as it directly captures the interruptions experienced throughout the day. The top correlation results for how different activities relate to users are displayed in Table 3. We also report the top predicting activities for each score and demographic group (i.e., male, female, ages 18-24, ages 25+).
As shown in Table 3, the correlations are weak with most activities, but few activities have moderate correlations. Tools related activities were positively correlated with overall valence and for both age groups. Particularly, Idle Tools have a relatively stronger correlation with valence in older adults, indicating an increase in the mood at work (Idle Tools include phone home screens and phone keyboards). Surfing by women and young age group are negatively correlated to valence, indicating reading news or surfing (e.g., reading COVID-19-related news) are negatively related to their valence. Productivity was negatively correlated with Communication Tools (e.g., keyboards and chat clients), and positively correlated with Business Communication (i.e., both Communication and Business apps such as Microsoft Teams). It can be seen that while both are Communication applications, whether or not they are used for work may imply productivity levels. Interruptions can be seen as the inverse of productivity, as it has positive correlation with Communication Tools, but the value is higher. The correlation reaches a moderate level for older adults while weakening for the young age group. This might be caused by the difference in working loads between young and older age groups. Lifestyle negatively correlates with interruptions, especially for female and older adults.
Table 3. Top correlations for all variables
@@ 5.3. App Usage Correlation
All Valence Work and Write (-0.07) Idle Tools (0.11)
Arousal Business (-0.05) Social Tools (0.15)
Productivity Communication Tools (-0.24) Business Communication (0.18)
Interruptions Lifestyle (-0.15) Communication Tools (0.29)
Male Valence Just Communication (-0.08) Business Communication (0.10)
Arousal Idling (-0.12) Just Communication (0.11)
Productivity Communication Tools (-0.15) Business Communication (0.34)
Interruptions Video Watching (-0.23) Communication Tools (0.18)
Female Valence Surfing (-0.17) Other (0.13)
Arousal Work and Write (-0.17) Social Tools (0.21)
Productivity Communication Tools (-0.30) Lifestyle (0.16)
Interruptions Lifestyle (-0.33) Idling (0.41)
Age 18-24 Valence Surfing (-0.13) Just Tools (0.09)
Arousal Idle Tools (-0.17) Social Tools (0.19)
Productivity Communication Tools (-0.21) Work and Write (0.20)
Interruptions Social Tools (-0.10) Communication Tools (0.16)
Age $\geq$25 Valence Just Communication (-0.12) Idle Tools (0.23)
Arousal Work and Write (-0.10) Idle Tools (0.15)
Productivity Idle Tools (-0.25) Social (0.20)
Interruptions Lifestyle (-0.24) Communication Tools (0.42)
Table 4. The computed features from smartphone usage behaviours
@@
sin_weekday Sine Transformation of weekday
@@
sin_time Sine Transformation of time period
cos_time Cosine Transformation of time period
@@
Notification num_noti Number of notifications
num_noti_interacted Number of notifications interacted
noti_cat Categorical Transformation of the number of notifications
@@
noti_interacted_pcg The percentage of notification interacted to the total number of notifications
@@
main_contact The main social relationship that contacted with
@@
main_activity The main physical activity of the users
num_activity Number of physical activities
@@
num_app The total number of apps
nunique_app_cat The number of unique app categories
app_noti The number of notification generated by non-communication apps
nunique_app_noti The number of unique apps sent notifications
main_cat_app The main category of apps
second_cat_app The second main category of apps
@@
6. Predicting User Emotion and Social Roles from Mobile Usage Data
In this section, we address RQ3 ”Can we model people’s emotion from mobile usage during home confinement? If so, what are the most significant types of features for prediction?”. First, we introduce the features extracted from mobile usage data and the prediction pipeline. We then list and discuss the prediction results of the general models and individual models. Finally, we examine the significant features for predicting human well-being.
@@ 6.1. Feature Extraction
Table 4 provides an overview of the features utilized in our research. Similar to previous mobile sensing studies (50; 81; 96), we extract contextual information, notifications-related information, physical activity, app, and activities inferred from apps. To process the app-usage data, we divide it into segments based on the ESM recording frequency (set at 90 minutes) and stored it in temporal sequences. Subsequently, we extract the features from these 90-minute segments. In order to handle cyclical features such as days (e.g., Monday) and times (e.g., Morning), we transform them into sine and cosine values. Additionally, it is worth noting that app usage patterns exhibit variations between weekdays and weekends (44; 50), prompting us to conduct separate analyses for these two categories. Regarding the time period per day, we divide it into six 4-hour segments, i.e., Early Morning, Morning, Afternoon, Evening, Night, and Late Night. Certain frequency-based features, such as notification-related or physical-activity-related features, are converted using quartiles, specifically the minimum, 25th percentile, 50th percentile, and 75th percentile. For instance, the number of interacted notifications is categorized into groups as ‘0’, ‘1-3’, ‘4-9’, and ‘10+’. Furthermore, the physical activities considered in our study only include active physical activities like bicycling, walking, and running.
To capture user activities related to app usage, we employed our proposed method, AIGTS, as detailed in Section 5.2. AIGTS allows us to extract both activities and their corresponding durations. Based on the extracted information, we compute the feature ‘entropy_user_activity’ by determining the proportion of each activity’s time portion within the entire 90-minute time interval. In addition to extracting individual features such as ‘num_noti’, we also generate composite features by combining two existing features. For example, the feature ‘num_communication’ is calculated by selecting and combining the notification sources. This composite feature counts the number of notifications originating the Business, Communication, or Social app categories. Notifications that do not belong to these app categories are counted as ‘app_noti’.
@@ 6.2. Prediction Pipeline
while emotion prediction can be addressed as a classification problem, with emotion levels categorized into two or three classes based on predefined thresholds (70; 98). We believe that utilizing regression is a better approach because valence and arousal labels are inherently ordinal data, where the relative order holds significance. In this research, we establish a regression-based pipeline to forecast the emotion scores of participants, as detailed below.
@@
Emotion. The ESM questionnaires, completed by the participants every 90 minutes, serve as the ground truth in the regression models developed in this study. Specifically, the item ‘valence’ was rated on a 5-point Likert scale from ‘unhappy’ to ‘happy’, where 1 indicates ‘unhappy’ and 5 refers to ‘happy’. Similarly, another item ‘arousal’ was also rated with a 5-point Likert scale from 1-5, with 1 indicating ‘calm’ and 5 means ‘excited’. To predict the levels of valence and arousal, the app activities within 90 minutes prior to completing the ESM questionnaire were leveraged as features.
@@
Regressors. In this research, we aim to develop regression models that accurately predicts human emotions during home confinement. To accomplish this, we employ three commonly used regression models: K-Nearest Neighbor (KNN) (84) regressor, Random Forest (RF) (83) regressor, and Gradient Boosting (GB) (29) regressor. KNN is a widely utilized model known for its versatility across various domains. It works by finding the K nearest neighbours to a given data point and using their values to predict the target variable. RF and GB, on the other hand, are ensemble models based on decision trees. These models combine predictions from multiple trees. However, there is a distinction in their approach. GB constructs trees or aggregates results in a way that minimizes the loss function, whereas RF constructs individual trees or aggregates results simultaneously.
@@
Feature Selection. Various methods, such as SHAP (54), LIME (77), and the $F$-statistic ($F$) (72), can be used for feature selection. In this study, we employ the $F$-statistic to determine the appropriate number of features. While most of our features are numeric, the $F$-statistic measures the univariate linear dependence between each feature and the target variable, taking into account statistical significance. Generally, a higher $F$ value indicates greater importance of the feature in predicting the target variable. In our experiments, we compare regression models using three feature sets: all features, features selected based on an $F$ value greater than or equal to 5, and features selected based on an $F$ value greater than or equal to 10.
@@
Validation. To compensate for the limited richness of our data, we have implemented nested cross-validation (CV) (61) to train and evaluate our prediction model. This approach allows us to accurately estimate the unbiased generalization performance of the model. The nested CV consists of an inner loop and an outer loop. Initially, the dataset is divided into training, validation, and testing sets. This helps optimize the model’s performance. In the outer loop, similar to previous human-centred studies (33; 25), we divide the data into multiple groups, each group representing one participant. To ensure unbiased evaluation, we apply k-fold cross validation (93) where k=5 on all groups. This means that the data from the same participants are not used in both the training and test sets simultaneously. In the inner loop, the remaining data groups are further split into three folds, with each fold serving as a validation set. We optimize the hyperparameters on the training set, evaluate them on the validation set, and select the best parameter settings based on the results from the three folds.
(a) MAE for valence
(b) MAE for arousal
(c) RMSE for valence
(d) RMSE for arousal
Figure 8. Prediction performance for valance and arousal with different feature sets, where the error bar presents standard deviation.
@@
Baselines. Similar to previous human-centered studies (33; 92), we compare the proposed model with two baselines. The first baseline is the Mean baseline, which calculates the average score of user responses. The second baseline is the Median baseline, which calculates the median value based on the data distribution and treats it as a predicted value. To evaluate the performance of regression models, we employ two commonly used metrics: Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) (19). Generally, lower values of MAE or RMSE indicate better prediction performance.

### v3

@@ 4.3.2. Unsupervised Activity Recognition From App Sequences
Segmentation. To investigate the relationship between app usage and various factors, we propose App-based Information Gain Temporal Segmentation (AIGTS), a method that focuses on identifying different activities within app usage. Figure 5 illustrates how activities are extracted from app usage data. AIGTS is inspired by Information Gain Temporal Segmentation (IGTS) (68), an unsupervised method for segmenting multivariate time series. IGTS identifies splits that minimize the weighted entropy across the multivariate time series, where each variable corresponds to whether a specific app category was used.
@@
Unlike IGTS, which focuses solely on the time series, AIGTS detects activity transitions, such as switching from Email or Browser usage (both in the Communication category) to browsing Reddit (in the News and Media category). This method can segment app usage data at various granularities, ranging from half a day to just 30 seconds. The level of segmentation is determined by the variable $k$, which represents the number of splits made in the time series. To ensure consistency in segment granularity across days, we compute the optimal $k$ using the following steps: (1) Identify the knee point of the activity data for each day; (2) Calculate the average number of segments per minute for all days based on the knee point; (3) Compute the median segments per minute across all users; (4) Multiply this median by the number of minutes each user spent on their device per day and select this value as $k$. These steps ensure that the average segment length remains consistent across all days.
@@
Figure 5. An illustration of unsupervised activity recognition using app usage data. The Segment Vector indicates the app subcategory.
@@
Clustering. Next, AIGTS is applied to the all-time series, resulting in a segmented list for each user. To extract features from these segments, K-means clustering (41) was applied. The distribution of different categories appearing in a segment was treated as an $n$ dimensional vector, where $n$ is the number of app categories. To identify the optimal number of segments, Silhouette Analysis (42) was utilized with a maximum of 20 clusters. Especially, the label indicating which cluster a segment belongs to was used as a feature representing the activity within that segment. These segment labels were derived manually from the average distribution of apps in that cluster. In this paper, a segment that belongs to a particular cluster will be referred to as an ‘activity’.
Figure 6. Proportion of app subcategory usage across user activity clusters. Clusters were manually named to describe the distribution of average activities within each cluster.
Overview of User Activities. Following segmentation and clustering, we identified 16 distinct clusters representing user activities, with an average activity duration of 18 minutes. Figure 6 presents a heatmap showing the proportion of app subcategories in each cluster. The $x$-axis denotes the app categories (categories that infrequently appear in each segment are omitted), while the $y$-axis represents the clusters (manually named such as Deep Work, Hybrid Work). The color intensity indicates the relative contribution of each app category to the cluster’s overall activity distribution.
The Device Idle cluster, corresponding to Personalization apps (e.g., launchers, as shown in Table 3), represents periods when users remain on their device’s home screen. We also found that the clusters Communication Only and Device Idle were the most frequent user activities during the COVID-19 home confinement, accounting for 17% and 14.6% of all activities, respectively. These activities are characterized by high proportions of communication-related apps and personalization apps, suggesting users spent considerable time on their devices in passive states or communicating with others. In addition, clusters such as Work-Related Communication and Hybrid Work exhibited higher proportions of productivity and business apps. These clusters likely reflect professional tasks and remote work activities, which were common during the lockdown period. The differentiation between Light Work and Deep Work clusters highlights variations in the intensity and type of work-related app usage.
4.3.3. Correlation Analysis of User Activity/App Usage and Well-Being
. We further investigate how user activities and app usage behaviours correlate with four key dimensions of well-being: valence, arousal, productivity, and interruptions. To assess these correlations, we performed Spearman Rank Correlation (18) tests between the time spent on different activities/apps and the scores for each dimension. Specifically, valence and arousal were examined for theirr relationship to the time spent on user activities in the preceding hour, while productivity and interruptions were calculated based on the end-of-day surveys.
Table 4. Top correlations between app usage/user acitivty and well-being dimensions.
@@ 4.3.3. Correlation Analysis of User Activity/App Usage and Well-Being
All Valence Hybrid Work (-0.07) Tools Apps (0.11)
Arousal Business Tasks (-0.05) Social Apps (0.15)
Productivity Communication Apps (-0.24) Work-Related Communication (0.18)
Interruptions Personal Tasks (-0.15) Communication Apps (0.29)
Male Valence Communication Only(-0.08) Work-Related Communication (0.10)
Arousal Device Idle (-0.12) Communication Only (0.11)
Productivity Communication Apps (-0.15) Work-Related Communication (0.34)
Interruptions Video Streaming (-0.23) Communication Apps (0.18)
Female Valence Surfing (-0.17) Miscellaneous (0.13)
Arousal Hybrid Work (-0.17) Social Apps (0.21)
Productivity Communication Apps (-0.30) Personal Tasks (0.16)
Interruptions Personal Tasks (-0.33) Device Idle (0.41)
Age 18-24 Valence Surfing (-0.13) Tools Apps (0.09)
Arousal Tools Apps (-0.17) Social Apps (0.19)
Productivity Communication Apps (-0.21) Hybrid Work (0.20)
Interruptions Social Apps (-0.10) Communication Apps (0.16)
Age $\geq$25 Valence Communication Only (-0.12) Tools Apps (0.23)
Arousal Hybrid Work (-0.10) Tools Apps (0.15)
Productivity Tools Apps (-0.25) Social (0.20)
Interruptions Personal Tasks (-0.24) Communication Apps (0.42)
Interestingly, as shown in Table 4, we found mostly weak correlations, with a few activities exhibiting moderate relationships. For instance, activities related to Tools Apps (e.g., phone home screens and keyboards) showed a moderate positive correlation with valence, especially among older adults. This suggests that certain habitual or passive tasks, such as interacting with productivity tools, may positively influence mood, possibly linked to work-related or routine activities. Additionally, Surfing activity (e.g., browsing news, including COVID-19 updates) showed a negative correlation with valence among women and the younger age group, indicating that news consumption during this period may be linked to lower mood. The correlation with productivity was notably negative for Communication Apps (e.g., keyboards, chat apps), but positively correlated with Work-Related Communication activity (e.g., business apps like Microsoft Teams), highlighting that whether a communication tool is used for professional or personal purposes may influence productivity levels.
Similarly, interruptions were mostly positively correlated with Communication Apps, particularly among older adults. This reflects the potentially disruptive nature of non-work-related communication during the day, which may vary by age. Conversely, Personal Tasks(e.g., personal apps or leisure activities) were negatively correlated with interruptions, especially for older adults and women, suggesting that non-work-related tasks may help buffer against disruptions, possibly offering a sense of control or relaxation.
5. Predicting User Emotion and Social Roles from Mobile Usage Data (RQ2)
In this section, we address RQ2 ”Can mobile usage data be used to model emotional well-being during home confinement, and which features are most predictive?”. First, we introduce the features extracted from mobile usage data and the prediction pipeline. We then list and discuss the prediction results of the general models and individual models. Finally, we examine the significant features for predicting human well-being.
Table 5. The computed features from smartphone usage behaviours. For the valence model, the features selected by F-statistic $\geq 5^{*}$ or $\geq 10^{**}$. For the arousal model, the features selected by F-statistic $\geq 5^{+}$ or $\geq 10^{++}$.
@@ 5. Predicting User Emotion and Social Roles from Mobile Usage Data (RQ2)
sin_weekday∗∗,+ Sine Transformation of weekday
@@
sin_time∗∗,++ Sine Transformation of time period
cos_time∗∗,++ Cosine Transformation of time period
@@
Notification num_noti∗∗,++ Number of notifications
num_noti_interacted∗∗,++ Number of notifications interacted
noti_cat++ Categorical Transformation of the number of notifications
@@
noti_interacted_pcg∗∗,++ The percentage of notification interacted to the total number of notifications
@@
main_contact∗∗ The main social relationship that contacted with
@@
main_activity∗∗,++ The main physical activity of the users
num_activity∗,+ Number of physical activities
@@
num_app∗∗,++ The total number of apps
nunique_app_cat∗∗ The number of unique app categories
app_noti∗∗,++ The number of notification generated by non-communication apps
nunique_app_noti∗∗,++ The number of unique apps sent notifications
main_cat_app∗,++ The main category of apps
second_cat_app∗∗,++ The second main category of apps
@@ 5.1. Feature Extraction
Table 5 provides an overview of the features utilized in our research. Similar to previous mobile sensing studies (40; 69; 86), we extract contextual information, notifications-related information, physical activity, app, and activities inferred from apps. To process the app-usage data, we divide it into segments based on the ESM recording frequency (set at 90 minutes) and stored it in temporal sequences. Subsequently, we extract the features from these 90-minute segments. In order to handle cyclical features such as days (e.g., Monday) and times (e.g., Morning), we transform them into sine and cosine values. Additionally, it is worth noting that app usage patterns exhibit variations between weekdays and weekends (35; 40), prompting us to conduct separate analyses for these two categories. Regarding the time period per day, we divide it into six 4-hour segments, i.e., Early Morning, Morning, Afternoon, Evening, Night, and Late Night. Certain frequency-based features, such as notification-related or physical-activity-related features, are converted using quartiles, specifically the minimum, 25th percentile, 50th percentile, and 75th percentile. For instance, the number of interacted notifications is categorized into groups as ‘0’, ‘1-3’, ‘4-9’, and ‘10+’. Furthermore, the physical activities considered in our study only include active physical activities like bicycling, walking, and running.
To capture user activities related to app usage, we employed our proposed method, AIGTS, as detailed in Section 4.3.2. AIGTS allows us to extract both activities and their corresponding durations. Based on the extracted information, we compute the feature ‘entropy_user_activity’ by determining the proportion of each activity’s time portion within the entire 90-minute time interval. In addition to extracting individual features such as ‘num_noti’, we also generate composite features by combining two existing features. For example, the feature ‘num_communication’ is calculated by selecting and combining the notification sources. This composite feature counts the number of notifications originating the Business, Communication, or Social app categories. Notifications that do not belong to these app categories are counted as ‘app_noti’.
@@ 5.2. Prediction Pipeline
While emotion prediction can be addressed as a classification problem, with emotion levels categorized into two or three classes based on predefined thresholds (58; 88). We believe that utilizing regression is a better approach because valence and arousal labels are inherently ordinal data, where the relative order holds significance. In this research, we establish a regression-based pipeline to forecast the emotion scores of participants, as detailed below.
Emotion
@@
The ESM questionnaires, completed by the participants every 90 minutes, serve as the ground truth in the regression models developed in this study. Specifically, the item ‘valence’ was rated on a 5-point Likert scale from ‘unhappy’ to ‘happy’, where 1 indicates ‘unhappy’ and 5 refers to ‘happy’. Similarly, another item ‘arousal’ was also rated with a 5-point Likert scale from 1-5, with 1 indicating ‘calm’ and 5 meaning ‘excited’. To predict the levels of valence and arousal, the app activities within 90 minutes prior to completing the ESM questionnaire were leveraged as features.
Regressors
@@
In this project, we aim to develop regression models that accurately predict human emotions during home confinement. To accomplish this, we employ three commonly used regression models: K-Nearest Neighbor (KNN) (72), Random Forest (RF) (71), and Gradient Boosting (GB) (21). KNN is a widely utilized model known for its versatility across various domains. It works by finding the K nearest neighbours to a given data point and using their values to predict the target variable. RF and GB, on the other hand, are ensemble models based on decision trees. These models combine predictions from multiple trees. However, there is a distinction in their approach. GB constructs trees or aggregates results in a way that minimizes the loss function, whereas RF constructs individual trees or aggregates results simultaneously.
To explore predictive capacity more comprehensively, we first develop general models using data from all participants combined. Next, recognizing that participants with similar routines or lifestyles may exhibit comparable patterns, we construct group models by dividing participants based on their self-reported habits. Finally, we analyze individual models to account for the unique characteristics and patterns exhibited by different individuals.
Feature Selection
@@
Various methods, such as SHAP (44), LIME (65), and the $F$-statistic ($F$) (60), can be used for feature selection. In this study, we employ the $F$-statistic to determine the appropriate number of features. While most of our features are numeric, the $F$-statistic measures the univariate linear dependence between each feature and the target variable, taking into account statistical significance. Generally, a higher $F$ value indicates greater importance of the feature in predicting the target variable. In our experiments, we compare regression models using three feature sets: all features, features selected based on an $F$ value greater than or equal to 5, and greater than or equal to 10.
Validation
@@
To address the limited richness of our data, we implemented Nested Cross-Validation (CV) (50) to train and evaluate our prediction models. Nested CV comprises an outer and an inner loop, both employing k-fold cross-validation (83) with different values of $k$. The outer loop splits the data into $k$ folds, using each fold for testing. The inner loop further divides the remaining data into $k$ folds, with each fold serving as a validation set to train the model and optimize hyperparameters. The best parameter settings are selected based on the inner loop results. This approach provides an accurate estimate of the model’s unbiased generalization performance.
For the general models, we use a 3-fold CV in the inner loop and a 5-fold CV in the outer loop, repeated over 3 trials. The group models employ 2-fold CV in the inner loop and 3-fold CV in the outer loop. Given the limited data available for individual participants, the individual models use a simpler 5-fold non-nested CV.
@@
Baselines and Metrics. Similar to previous human-centered studies (24; 82), we compare the proposed models with two baselines, Mean, which calculates the average score of user responses, and Median, which calculates the median value based on the data distribution and treats it as a predicted value. To evaluate the performance of regression models, we employ two commonly used metrics: Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) (13). Generally, lower values of MAE or RMSE indicate better prediction performance.